# Partial Mobilization: Tracking Multilingual Information Flows Amongst Russian Media Outlets and Telegram

### v4

@@ 1 Introduction
We find that much of the content shared between Telegram and Russian websites concerns the war in Ukraine and Western sanctions on Russia. By performing this same clustering on semantic content specific to only Russian news websites or only Telegram channels, we find an emphasis on the day-to-day machinations (e.g., bombings of particular bridges) of the Russo-Ukrainian War on Telegram that contrasts with a focus on US politics specific on Russian news outlets. Next, we track the spread of topics amongst and between Telegram and Russian news media. We find that 33.2% of distinct topics discussed on our set of Telegram channels, making up 24.3% of all Telegram messages in our dataset, were first published in Russian news articles. In contrast, 13.9% of topics on Russian news websites, comprising 18.39% of all the text (all messages from Telegram and paragraphs from Russian state media) on our set of Russian websites, were posted first on Telegram. Telegram-originated messages comprise a particularly large amount of content on news websites like waronfakes.com (28.2% of text), ukraina.ru (27.9%), and ura.news (25.6%) that all maintain large Telegram presences. Applying time-series analysis with the Hawkes processes on our topic clusters, we then estimate the percentage of content on each platform that was influenced by activity on other platforms. While some websites like ura.news had relatively low amounts of content (2.4%) flowing from Telegram, others such as ukraina.ru had much larger (27.2%) percentages of their content possibly influenced by activity on Telegram. Finally, we analyze how quickly topics spread amongst our set of news websites and onto Telegram channels, finding that websites like ura.news, ukraina.ru, and news-front.info, and the Telegram channel @gensham’s topics flow most quickly to other platforms.
@@ 2 Background and Related Work
Telegram. Telegram, started in 2013, is a messaging platform  (Baumgartner et al. 2020) with 700 million monthly users (Singh 2022) as of June 2022. Similar to WhatsApp and Facebook Messenger, users on Telegram can share messages, images, and videos. However, in addition to private conversations, Telegram users can create one-to-many public channels (where only channel creators and administrators can post content) to which other users can subscribe (La Morgia et al. 2021). Within this work, we focus on these administrator-run channels.
@@
Platform Art. Embed.
Telegram 2,477,564
geopolitca.ru 544 4,223
globalresearch.ca 5,904 158,213
govoritmoskva.ru 57,887 223,887
journal-neo.org 2,509 15,363
katehon.com 4,393 45,832
lug-info.com 1,221 89,27
news-front.info 29,052 261,931
rbc.ru 15,994 179,273
Platform Art. Embed.
rt.com 12,701 138,252
southfront.org 7,785 78,963
sputniknews.com 18,214 183,218
strategic-culture.org 2,174 25,096
tass.com 9,106 44,877
ukraina.ru 17,435 193,562
ura.news 7,919 125,619
waronfakes.com 2,478 10,890
@@ 4 Methodology
We compare the semantic content of our embedded messages and paragraphs utilizing cosine similarity (Song et al. 2020). As found in previous works (Hanley, Kumar, and Durumeric 2023b; Vetter et al. 2022; Song et al. 2020; Bernard et al. 2022), a cosine similarity threshold between 0.60 and 0.80 can be utilized to determine whether two pieces of text are about the same topic. For instance, with the same model, Phan et al. (2022), found that a threshold near 0.715 achieved the best results. However, in order to further verify these past results, we perform our own evaluation to determine a threshold at which two messages/paragraphs can reliably be said to be about the same topic.
We take two approaches to determine the appropriate threshold for considering two paragraphs to be about the same topic. First, we take 250 random paragraph pairs with similarities at various thresholds (i.e, for 0.60, messages/paragraphs with similarities between 0.59 and 0.61) and have an expert determine whether they are about the same topic as outlined in Hanley, Kumar, and Durumeric (2023b); Soper et al. (2021). We perform this evaluation in monolingual settings for English and Russian (the two most prominent languages in our dataset) and a multilingual setting with English and Russian. Given that our expert could not speak Russian, we utilized Google Translate to translate our set of Russian paragraphs into English. As seen in Table 2, our selected pre-trained model achieves a near 85.2% topic-similarity precision at a threshold of 0.7 only for English. For Russian, and in a multilingual setting of English and Russian, our model only achieves this precision at a threshold of 0.8. This is in line with prior work (Verma et al. 2022) that has found that multilingual models often over-perform in English while underperforming in rarer languages.
@@
Obrázek 1: Following, the Russian invasion of Ukraine, the number of Telegram posts in our set of 732 channels spiked from an average of 6,500 to over 27,500. Similarly, our set of Russian websites began to utilize Western media sites like Twitter less often while increasing their use of Telegram, with Telegram increasing from 4% of their external hyperlinks to nearly 10%. Our set of Telegram channels themselves began to steadily use platforms like YouTube less often.
@@ 5 Behavior of Russian State Media and Associated Telegram Channels
Website Most Similar Telegrams Similarities
geopolitika.ru rossiyaneevropa, pintofmind, russtrat 0.116, 0.099, 0.090
globalresearch.ca gr_crg, karaulny, oleg_blokhin 0.109, 0.105, 0.089
govoritmoskva.ru radiogovoritmsk, karaulny, vzglyad_ru 0.340, 0.301, 0.263
journal-neo.org rossiyaneevropa, pintofmind, russtrat 0.095, 0.081, 0.074
katehon.com rossiyaneevropa, vzglyad_ru, rus_demiurge 0.364, 0.362, 0.354
lug-info.com lic_lpr,gtrklnr_lugansk24,dnr_sckk 0.516, 0.332, 0.313
news-front.info ukraina_ru, vzglyad_ru, rossiyaneevropa 0.473, 0.458, 0.444
rbc.ru rbc_news, radiogovoritmsk, vzglyad_ru 0.486, 0.433, 0.415
rt.com vzglyad_ru, radiogovoritmsk, rbc_news 0.135, 0.121, 0.101
southfront.org southfronteng,bloknot_rossii, nm_dnr 0.150, 0.125, 0.108
sputniknews.com vzglyad_ru, radiogovoritmsk, tass_agency 0.125, 0.106, 0.101
strategic-culture.org strategic_culture, rossiyaneevropa, pintofmind 0.087, 0.084, 0.079
tass.com tassagency_en, tass_agency, vzglyad_ru 0.273, 0.234, 0.193
ukraina.ru ukraina_ru, vzglyad_ru, rossiyaneevropa 0.526, 0.495, 0.483
ura.news vzglyad_ru, wek_ru, leningrad_guide 0.354, 0.311, 0.311
waronfakes.com warfakes, waronfakesen, warfakebelgorod 0.299, 0.277, 0.170
Tabulka 4: Most similar Telegram chanels for each website.
Obrázek 3: Propaganda image from @rossiyaneevropa. In the post, @rossiyaneevropa argues that NATO cannot criticize Russia’s activities in Ukraine given the West’s war crimes in Libya.
@@ 6 The Spread of Topics
% of Content First % of Content First
@@
geopolitika.ru rossiyaneevropa, karaulny, uraldaily 0.47%, 0.45%, 0.38%
globalresearch.ca karaulny, traugottickerothliveticker, infantmilitario 0.32%, 0.29, 0.16%
govoritmoskva.ru karaulny, krimski, vibornyk 1.5%, 0.61%, 0.49%
journal-neo.org parstodayrussian, karaulny, vsya_korea 0.50%, 0.42%, 0.41%
katehon.com karaulny,new_militarycolumnist, marochkolive 1.0%, 0.75%, 0.71%
lug-info.com lic_lpr,donbassr,mid_lnr 2.7%, 1.2%, 0.72%
news-front.info karaulny, new_militarycolumnist, solovievlive 1.2%, 1.1%, 0.69%
rbc.ru karaulny, bbbreaking, infantmilitario 1.6%, 0.60%, 0.50%
rt.com karaulny, sportrian, infantmilitario 0.43%, 0.39%, 0.29%
southfront.org new_militarycolumnist,karaulny, intelslava 0.57%, 0.42%, 0.40%
sputniknews.com karaulny, infantmilitario, parstodayrussian 0.38%, 0.31%, 0.28%
strategic-culture.org karaulny, vityzeva, russtrat 0.37%, 0.21%, 0.21%
tass.com karaulny, ontnews, bbbreaking 0.61%, 0.45%, 0.41%
ukraina.ru karaulny, solovievlive, donbassr 1.7%, 1.1%, 0.72%
ura.news karaulny, russica2, donbassr 2.6%, 1.7%, 1.2%
waronfakes.com senkevichonline, mykolaivskaoda, dnronline 5.5%, 5.5%, 5.1%
Tabulka 7: Top three telegrams — by percentage of content first posted — that post content prior to our set of websites.
@@
Examining the set of Telegram channels that most often first posted each website’s content in Table 7, we see the Telegram channels @kalaruny (159K subscribers), @new_militarycolumnist (213K), and @bbbreaking (1.38M) often posted content first across nearly all of our websites. As reported elsewhere, @karaulny (screen name Karaulny-Z) is a pro-Russian government channel that abides by a “stop list” of prohibited topics and that was purchased by supporters of the Kremlin in 2017 (Rothrock 2018). @new_militarycolumnist, another pro-Russian channel, gives continuous updates on the Russian military (Sabah 2020). Finally, @bbbreaking, another pro-Russian Russian-language channel, with the motto “Earlier than others. Almost” appears to also in several cases give updates on the news before many of our Russian news websites.
@@
Plotting the estimated influence of each website on each other and Telegram, we see in Figure 4, that Telegram typically has a moderate estimated influence on the content published on each of the websites. Most prominently, Telegram has a possible influence on 18.4% of content on katehon.com, 25.2% on news-front.info, 23.0% on rbc.ru, and 26.7% on ukraina.ru. Conversely, it has the weakest influence on the website ura.news (despite often writing about its topics prior to ura.news [Table 6]). Conversely, with the smallest amount of articles, we see that geopolitika.ru also had the smallest estimated influence on Telegram. Further as largely expected, we see in the estimated topic spread efficiency in Figure 5 that many of our websites are somewhat efficient at seeing their content migrate to Telegram, with waronfakes.com being the best at 18.8% efficiency. This indicates across our websites, many may have to post a few articles before that article’s topic appears on Telegram.
@@
Spread of Topics From Telegram Channels. We now examine the rate at which topics spread from some Telegram channels amongst themselves and to Russian websites. We display the set of 16 Telegram channels whose content spreads the furthest (i.e., the channels that originated topics that then spread prolifically and widely) in Figure 7. As seen in Figure 7, the Telegram channel most effective at originating content/topics that are then reposted elsewhere both on Telegram and within the Russian website ecosystem is @gensab (86.9K subscribers). A pro-Russian propaganda channel, the channel continuously comments on the Russian invasion of Ukraine, writing on August 29th:
@@ 7 Discussion and Conclusion
Identifying Propaganda Channels. As briefly discussed in Section 5, our approach can also be used to identify new Telegram channels that align with Russian state-promoted narratives. As seen in Table 4, our method identified Telegram channels that supported anti-Ukrainian and pro-Russian-separatist sentiments in an automated fashion (e.g., @lic_lpr, @gtrklnr_lugansk24, @dnr_sckk) as well as several channels that largely repeated Russian state-media topics (e.g., @rossiyaneevropa, @pintofmind, @russtrat) (Afroz and Sehgal 2022). We note that this method can further be utilized beyond Russian and Ukrainian-focused topics to identify websites that are similar to disinformation, hyperpartisan, or other types of malicious websites.
Content Spreading From Different Platforms. By utilizing MPNet, DP-Means clustering, and Hawkes processes, we estimated the role of Telegram and identified the most influential Russian websites within our ecosystem. Unlike past work that has relied on the presence of hyperlinks (Hanley, Kumar, and Durumeric 2022; La Morgia et al. 2021), our approach approximates influence by directly measuring topics themselves. This approach can be extended to understand how larger platforms like Facebook, Reddit, and Twitter interact with Russian propaganda. While we limited ourselves to analyzing the semantic correspondence of Russian websites, our approach can be used to estimate the influence of news websites more generally (e.g., cnn.com, foxnews.com).
@@
  • La Morgia et al. (2021) La Morgia, M.; Mei, A.; Mongardini, A. M.; and Wu, J. 2021. Uncovering the Dark Side of Telegram: Fakes, Clones, Scams, and Conspiracy Movements. arXiv preprint arXiv:2111.13530.
@@
  • Vetter et al. (2022) Vetter, D.; Tithi, J. J.; Westerlund, M.; Zicari, R. V.; and Roig, G. 2022. Using Sentence Embeddings and Semantic Similarity for Seeking Consensus when Assessing Trustworthy AI. arXiv:2208.04608.

### v5

# creation2024-04-01T00:00:00-05:00

@@ 1 Introduction
We find that much of the content shared between Telegram and Russian websites concerns the war in Ukraine and Western sanctions on Russia. By performing this same clustering on semantic content specific to only Russian news websites or only Telegram channels, we find an emphasis on the day-to-day machinations (e.g., bombings of particular bridges) of the Russo-Ukrainian War on Telegram that contrasts with a focus on US politics specific on Russian news outlets. Next, we track the spread of topics amongst and between Telegram and Russian news media. We find that 33.2% of distinct topics discussed on our set of Telegram channels, making up 24.3% of all Telegram messages in our dataset, were first published in Russian news articles. In contrast, 13.9% of topics on Russian news websites, comprising 18.4% of all the text (all messages from Telegram and paragraphs from Russian state media) on our set of Russian websites, were posted first on Telegram. Telegram-originated messages comprise a particularly large amount of content on news websites like waronfakes.com (28.2% of text), ukraina.ru (27.9%), and ura.news (25.6%) that all maintain large Telegram presences. Applying time-series analysis with the Hawkes processes on our topic clusters, we then estimate the percentage of content on each platform that was influenced by activity on other platforms. While some websites like ura.news had relatively low amounts of content (2.3%) flowing from Telegram, others such as ukraina.ru had much larger (26.7%) percentages of their content possibly influenced by activity on Telegram. Finally, we analyze how quickly topics spread amongst our set of news websites and onto Telegram channels, finding that websites like ura.news, ukraina.ru, and news-front.info, and the Telegram channel @genshab’s topics flow most quickly to other platforms.
@@ 2 Background and Related Work
Telegram. Telegram, started in 2013, is a messaging platform  (Baumgartner et al. 2020) with 700 million monthly users (Singh 2022) as of June 2022. Similar to WhatsApp and Facebook Messenger, users on Telegram can share messages, images, and videos. However, in addition to private conversations, Telegram users can create one-to-many public channels (where only channel creators and administrators can post content) to which other users can subscribe. Within this work, we focus on these administrator-run channels.
@@
Platform Art. Emb.
Telegram 2.48M
geopolitca.ru 544 4.22K
global-research.ca 5.90K 158K
govorit-moskva.ru 57.9K 224K
journal-neo.org 2.51K 15.4K
katehon.com 4.39K 45.8K
lug-info.com 1.22K 8.92K
news-front.info 29.0K 262K
Platform Art. Emb.
rbc.ru 16.0K 179K
rt.com 12.7K 138K
southfront.org 7.79K 79.0K
sputnik-news.com 18.2K 183K
strategic-culture.org 2.17K 25.1K
tass.com 9.11K 44.9K
ukraina.ru 17.4K 193.6K
ura.news 7.92K 125.6K
waronfakes.com 2.48K 10.9K
@@ 4 Methodology
We compare the semantic content of our embedded messages and paragraphs utilizing cosine similarity (Song et al. 2020). As found in previous works (Hanley, Kumar, and Durumeric 2023b; Song et al. 2020; Bernard et al. 2022), a cosine similarity threshold between 0.60 and 0.80 can be utilized to determine whether two pieces of text are about the same topic. For instance, with the same model, Phan et al. (2022), found that a threshold near 0.715 achieved the best results. However, in order to further verify these past results, we perform our own evaluation to determine a threshold at which two messages/paragraphs can reliably be said to be about the same topic.
We take two approaches to determine the appropriate threshold for considering two paragraphs to be about the same topic. First, we take 250 random paragraph pairs with similarities at various thresholds (i.e., for 0.60, messages/paragraphs with similarities between 0.59 and 0.61) and have an expert determine whether they are about the same topic as outlined in Hanley, Kumar, and Durumeric (2023b); Soper et al. (2021). We perform this evaluation in monolingual settings for English and Russian (the two most prominent languages in our dataset) and a multilingual setting with English and Russian. Given that our expert could not speak Russian, we utilized Google Translate to translate our set of Russian paragraphs into English. As seen in Table 2, our selected pre-trained model achieves a near 85.2% topic-similarity precision at a threshold of 0.7 only for English. For Russian, and in a multilingual setting of English and Russian, our model only achieves this precision at a threshold of 0.8. This is in line with prior work (Verma et al. 2022) that has found that multilingual models often over-perform in English while underperforming in rarer languages.
@@
Obrázek 1: Following, the Russian invasion of Ukraine, the number of Telegram posts in our set of 732 channels spiked from an average of 6,500 to over 27,500. Similarly, our set of Russian websites began to utilize Western media sites like Twitter less often while increasing their use of Telegram (Telegram links increased from 4% of these websites’ external hyperlinks to nearly 10%). Our set of Telegram channels themselves began to steadily use platforms like YouTube less often.
@@ 5 Behavior of Russian State Media and Associated Telegram Channels
Website Sim. Telegram Simalarity
geopolitika.ru rossiyaneevropa 0.116
globalresearch.ca gr-crg 0.109
govoritmoskva.ru radiogovoritmsk 0.340
journal-neo.org rossiyaneevropa 0.095
katehon.com rossiyaneevropa 0.364
lug-info.com lic-lpr 0.516
news-front.info ukraina_ru 0.473
rbc.ru rbc_news 0.486
rt.com vzglyad-ru 0.135
southfront.org southfronteng 0.150
sputniknews.com vzglyad_ru 0.125
strategic-culture.org strategic_culture 0.087
tass.com tassagency_en 0.273
ukraina.ru ukraina_ru 0.526
ura.news vzglyad_ru 0.354
waronfakes.com warfakes 0.299
Tabulka 4: Most similar Telegram chanels to each website.
Obrázek 3: Propaganda image from @rossiyaneevropa. @rossiyaneevropa argues NATO cannot criticize Russia’s activities in Ukraine given the West’s war crimes in Libya.
@@ 6 The Spread of Topics
% First % First
@@
geopolitika.ru rossiyaneevropa 0.47%
globalresearch.ca karaulny 0.32%
govoritmoskva.ru karaulny 1.5%
journal-neo.org parstodayrussian 0.50%
katehon.com karaulny 1.0%
lug-info.com lic_lpr 2.7%
news-front.info karaulny 1.2%
rbc.ru karaulny 1.6%
rt.com karaulny 0.43%
southfront.org new_militarycolumnist 0.57%
sputniknews.com karaulny 0.38%
strategic-culture.org karaulny 0.37%
tass.com karaulny 0.61%
midrule ukraina.ru karaulny 1.7%
ura.news karaulny 2.6%
waronfakes.com senkevichonline 5.5%
Tabulka 7: Top telegrams — by percentage of content first posted — that post content prior to our set of websites.
@@
Examining the set of Telegram channels that most often first posted each website’s content in Table 7, we see the Telegram channels @kalaruny (159K subscribers), @new_militarycolumnist (213K), and @bbbreaking (1.38M; not shown in table) often posted content first across nearly all of our websites. As reported elsewhere, @karaulny (screen name Karaulny-Z) is a pro-Russian government channel that abides by a “stop list” of prohibited topics and that was purchased by supporters of the Kremlin in 2017 (Rothrock 2018). @new_militarycolumnist, another pro-Russian channel, gives continuous updates on the Russian military (Sabah 2020). Finally, @bbbreaking, another pro-Russian Russian-language channel, with the motto “Earlier than others. Almost” appears to also in several cases give updates on the news before many of our Russian news websites.
@@
Plotting the estimated influence of each website on each other and Telegram, we see in Figure 4, that Telegram typically has a moderate estimated influence on the content published on each of the websites. Most prominently, Telegram has a possible influence on 18.4% of content on katehon.com, 25.2% on news-front.info, 23.0% on rbc.ru, and 26.7% on ukraina.ru. Conversely, it has the weakest influence on the website ura.news (despite often writing about its topics prior to ura.news [Table 6]). With the smallest amount of articles, we see that geopolitika.ru also had the smallest estimated influence on Telegram while news-front.info has greatest effect on Telegram. Further as largely expected, we see in the estimated topic spread efficiency in Figure 5 that many of our websites are somewhat efficient at seeing their content migrate to Telegram, with waronfakes.com being the best at 18.8% efficiency. This indicates across our websites, many may have to post a few articles before that article’s topic appears on Telegram.
@@
Spread of Topics From Telegram Channels. We now examine the rate at which topics spread from some Telegram channels amongst themselves and to Russian websites. We display the set of 16 Telegram channels whose content spreads the furthest (i.e., the channels that originated topics that then spread prolifically and widely) in Figure 7. As seen in Figure 7, one of the Telegram channel most effective at originating content/topics that are then reposted elsewhere both on Telegram and within the Russian website ecosystem is @genshab (86.9K subscribers). A pro-Russian propaganda channel, the channel continuously comments on the Russian invasion of Ukraine, writing on August 29th:
@@ 7 Discussion and Conclusion
Identifying Propaganda Channels. As briefly discussed in Section 5, our approach can also be used to identify new Telegram channels that align with Russian state-promoted narratives. As was seen in Table 4, our method identified Telegram channels that supported anti-Ukrainian and pro-Russian-separatist sentiments in an automated fashion (e.g., @lic_lpr, @gtrklnr_lugansk24, @dnr_sckk) as well as several channels that largely repeated Russian state-media topics (e.g., @rossiyaneevropa, @pintofmind, @russtrat) (Afroz and Sehgal 2022). We note that while this method requires lists of predefined propaganda or disinformation platforms/websites, our method can further be utilized beyond Russian and Ukrainian-focused topics to identify websites that are similar to disinformation, hyperpartisan, or other types of malicious websites.
Content Spreading From Different Platforms. By utilizing MPNet, DP-Means clustering, and Hawkes processes, we estimated the role of Telegram and identified the most influential Russian websites within our ecosystem. Unlike past work that has relied on the presence of hyperlinks (Hanley, Kumar, and Durumeric 2022), our approach approximates influence by directly measuring topics themselves. This approach can be extended to understand how larger platforms like Facebook, Reddit, and Twitter interact with Russian propaganda. While we limited ourselves to analyzing the semantic correspondence of Russian websites, our approach can be used to estimate the influence of news websites more generally (e.g., cnn.com, foxnews.com).